OpenStudy (andijo76):

In the week before and the week after a holiday there were 10,000 total deaths and 4968 of them occurred in the week before the holiday. a) construct a 90% confidence interval estimate of the proportion of deaths in the week before the holiday to the total death in the week before and the week after the holiday b) based on the result does there appear to be any indication that people can temporarily postpone their death to survive the holiday

OpenStudy (anonymous):

R.I.P Math

OpenStudy (andijo76):

i agree haha but i still have to do it

OpenStudy (anonymous):

CI = mean proportion ± (z*)sqrt[ (p)(1-p) / n] where z* = 1.645 with 10%/2 = 0.05 area in each tail mean proportion p = 4946 / 10000 = 0.4946 now fill in the values from the problem ... CI = mean proportion ± (z*)sqrt[ (p)(1-p) / n] CI = 0.4946 ± (1.645)sqrt[ (0.4946)(1- 0.4946) / 10000] CI = (0.4864, 0.5028) hope that helps

OpenStudy (andijo76):

sorta does thanks I am very new at this class actually all the math is i think if i would have had a taste of the harder math in highschool it would have helped

OpenStudy (anonymous):

haha yes I'm in the high school stats atm lol this one was hard even 4 me 0.0